# LLMs’ Reshaping of People, Processes, Products, and Society in Software Development: A Comprehensive Exploration with Early Adopters

### v1

@@ 6. Threats to validity
As with any qualitative study, our research also faces several threats to validity, including:
6.1. Internal Validity:
The study’s timeline coincided with frequent updates to tools such as GPT-3.5 and Bard (now Gemini), creating a significant internal validity threat. The rapid evolution of these tools may have influenced participants’ experiences in uncontrolled ways. As participants interacted with different versions of the tools, they might have encountered different performance levels, which could skew the results. Consequently, findings related to GPT-3.5 or earlier versions of Bard may not accurately reflect the current version of these tools, especially with the emergence of newer iterations like GPT-4o and OpenAI o1.
Additionally, the absence of pretests to assess participants’ baseline knowledge of how they used LLMs may impact the study’s internal validity. Although post-survey questions regarding industry experience partially mitigated this issue, the lack of a uniform measure of skill at the beginning of the study means that differences in participants’ performance could be attributed to their varying levels of expertise rather than the tools themselves.
6.2. External Validity:
@@
Moreover, the lack of diversity in the participant pool, predominantly composed of White or Asian males with over 3 years of experience, further exacerbates the threat to external validity. The experiences and challenges faced by underrepresented groups—including women, non-binary individuals, developers with disabilities, and those from various racial, ethnic, and experience backgrounds—may differ significantly from those of the current sample. Although our study was announced openly on LinkedIn, we recognize the need to implement additional strategies in future research to ensure a more diverse and representative participant pool.
6.3. Construct Validity:
The introduction of newer models, such as GPT-4o, along with the resolution of issues related to GPT-3.5 and Bard, poses a significant threat to the temporal relevance of the findings. As LLM tools continue to evolve, the capabilities and limitations of earlier versions may no longer accurately represent the current state of technology. Consequently, some results from the study could become outdated, potentially undermining the applicability and relevance of the findings to contemporary tools usage. This emphasizes the necessity of considering the dynamic nature of LLM tools when interpreting the study’s conclusions. Additionally, the limited number of participants using tools like Google Bard and GitHub Copilot Chat presents a construct validity threat. The conclusions drawn about these tools may lack robustness and comprehensiveness due to insufficient data.
6.4. Conclusion Validity:
The relatively small sample size of 16 participants and 16 hours of interviews introduces threat to conclusion validity. This limitation restricts the statistical power to draw meaningful generalizations from the findings. Additionally, the small dataset heightens the risk of random variations influencing the results, potentially leading to unreliable conclusions. As a result, the study’s ability to make robust claims about the broader population of developers and their experiences with LLM tools may be compromised.
7. Conclusion
@@
The sixteen interviewed developers show an advanced understanding of how LLMs work and their data sources. This understanding influences the decisions that developers make about when, how, and why to use LLMs. Their insights can be used to craft best practices for LLM use in computing education and workforce settings. In addition, they demonstrate that LLMs offer, in general, far more opportunities than they do challenges. Despite concerns over code quality and limitations, our participants deemed LLMs to be sufficiently useful for developers. Consequently, our findings suggest that LLMs can be a valuable asset in a professional developer’s toolbox. As one participant expressed, “[ChatGPT] is like calculators being invented. We’re going to ban them for a while, and we’re going to tell [students] no, no, no, don’t use it. And then, once they graduate, they will use it every day.
8. Acknowledgments
This material is based upon work supported by the Air Force Office of Scientific Research under award number FA9550-21-1-0108 and National Science Foundation (NSF) under award numbers IIS-2313890, CCF-2006977, and IIS-1917885. Any opinions, findings, and conclusions or recommendations expressed in this material are those of the authors and do not necessarily reflect the view of the AFOSR or NSF.

### v2

@@ 6 Threats to validity
As a qualitative study, our research faces several threats to validity that should be considered when interpreting the findings.
6.1 Internal Validity:
The study’s timeline overlapped with a period of rapid change in LLM tools (e.g., early versions of GPT-3.5 and Google Bard), which may have influenced participants’ experiences in ways we could not fully control. Because interviews were conducted over several months, different participants may have interacted with slightly different model versions or configurations, making it difficult to separate tool evolution from differences in individual experience, tasks, or organizational context. As a result, our findings should be interpreted as characterizing early LLM-assisted development in general, rather than the behavior of any specific model.
Additionally, we did not administer a pretest to systematically assess participants’ baseline proficiency with LLM tools. We partially addressed this by restricting our sample to full-time software professionals and post-survey questions about industry experience. However, without a uniform baseline measure of LLM-specific skill, some observed differences in practices or outcomes may reflect pre-existing expertise rather than effects of the tools themselves.
6.2 External Validity:
@@
Moreover, the lack of diversity in the participant pool, predominantly composed of White or Asian males with over 3 years of experience, further exacerbates the threat to external validity. The experiences and challenges faced by underrepresented groups, including women, non-binary individuals, developers with disabilities, and those from various racial, ethnic, and experience backgrounds, may differ significantly from those of the current sample. Although our study was announced openly on LinkedIn, we recognize the need to implement additional strategies in future research to ensure a more diverse and representative participant pool.
6.3 Construct Validity:
Our constructs, such as “productivity,” “learning,” and “impact on process,” are based on participants’ self-reports in interviews rather than direct observation of work practices or objective productivity metrics. As a result, our findings may be influenced by recall bias, social desirability, or participants’ own interpretations of these constructs. We attempted to mitigate this by asking for examples, and triangulating across multiple participants, but our conclusions still reflect perceived rather than directly measured effects. A second threat is the rapid evolution of LLM tools. Our results therefore speak most directly to LLM coding assistants in 2023, and should be interpreted as patterns of LLM-mediated work rather than properties of any specific current model. Finally, only a minority of participants used tools such as Google Bard or GitHub Copilot Chat, so tool-specific observations, especially about less-used systems, should be treated as tentative. Our primary focus is on cross-tool practices and perceptions, rather than comparative evaluation of individual products.
6.4 Conclusion Validity:
With 16 participants and approximately 16 hours of interview data, our study was designed for in-depth thematic analysis rather than statistical generalization. This sample size is consistent with qualitative recommendations for achieving thematic saturation in relatively homogeneous populations, but it nonetheless limits our ability to make strong claims about prevalence or effect sizes in the broader developer population. We mitigated threats to conclusion validity through standard qualitative techniques, including independent coding by multiple researchers, iterative refinement of the codebook, and consensus-building discussions. Theme counts (e.g., how many participants mentioned a given practice) should therefore be interpreted as descriptive signals within our sample, not as precise estimates of how common those practices are in industry at large.
7 Conclusion
@@
The sixteen interviewed developers show an advanced understanding of how LLMs work and their data sources. This understanding influences the decisions that developers make about when, how, and why to use LLMs. Their insights can be used to develop best practices for the use of LLMs in computing education and workforce settings. Since our interviews were conducted in 2023, we explicitly compared them with more recent studies, and found that early reports of productivity gains, security and privacy concerns, phase-dependent use across the SDLC, and prompt-engineering practices have largely persisted, but some initial expectations about code quality, test automation, and the impact on junior roles have been tempered. Overall, our participants viewed LLMs as offering more opportunities than challenges: despite concerns over code quality and limitations, our participants deemed LLMs to be sufficiently useful for developers. Consequently, our findings suggest that LLMs can be a valuable asset in a professional developer’s toolbox. As one participant expressed, “[ChatGPT] is like calculators being invented. We’re going to ban them for a while, and we’re going to tell [students] no, no, no, don’t use it. And then, once they graduate, they will use it every day.
Declarations
Funding
This material is based upon work supported by the Air Force Office of Scientific Research under award number FA9550-21-1-0108 and the National Science Foundation (NSF) under award numbers IIS-2313890, CCF-2006977, and IIS-1917885. Any opinions, findings, and conclusions or recommendations expressed in this material are those of the authors and do not necessarily reflect the view of the AFOSR or NSF.
Ethical approval
This study was reviewed and approved by the North Carolina State University Institutional Review Board (Protocol #25913).
Informed consent
All participants provided informed consent prior to participation, including consent for anonymized quotations to be published.
Author contributions
Benyamin Tabarsi and Heidi Reichert contributed equally to this work. Material preparation and data collection were led by Benyamin Tabarsi and Heidi Reichert. Initial coding and analysis were conducted by Benyamin Tabarsi and Heidi Reichert, with additional support from Ally Limke during data tagging and Sandeep Kuttal during tag grouping. The first draft of the manuscript was written by Benyamin Tabarsi and Heidi Reichert. Sam Gilson contributed substantially to refining and updating multiple sections of the manuscript, particularly the related work and discussion sections. Sandeep Kuttal and Tiffany Barnes served as supervising authors, providing guidance on study design, analysis, and interpretation of results. All authors (Benyamin Tabarsi, Heidi Reichert, Sam Gilson, Ally Limke, Sandeep Kuttal, and Tiffany Barnes) reviewed, commented on, and approved the final manuscript.
Data availability
The interview data that support the findings of this study contain information that could compromise participant privacy and are therefore subject to Institutional Review Board (IRB) restrictions (NC State – Protocol #25913). In keeping with these ethical requirements, full audio recordings and verbatim transcripts are not publicly available.
Conflict of interest
The authors have no relevant financial or non-financial interests to disclose.
Clinical trial number
Clinical trial number: not applicable.